OpenStudy (anonymous):

Joe purchased some blue suede material to use as the surface of his pond. The amount of suede used to cover the pond cost about $13.10. What is a good estimate of the cost of the suede to cover the real-life pond?

OpenStudy (anonymous):

need specifics

OpenStudy (anonymous):

the ratio is 1:12

OpenStudy (anonymous):

that website has the problems on it

OpenStudy (anonymous):

Just multiply by 12, so \[13.10*12=157.2\]

OpenStudy (anonymous):

Pretty sure that is correct, and i cant access your doc. it says Bad request

OpenStudy (anonymous):

you need to copy and paste the whole thing

OpenStudy (anonymous):

Yeah my answer is correct,

OpenStudy (anonymous):

for number 5. would the answer be 1.71 feet?

OpenStudy (anonymous):

Yeah, thats what I'm getting well done !

OpenStudy (anonymous):

will you help me with 6 and 9 also please

OpenStudy (anonymous):

\[12*7= 84\]

OpenStudy (anonymous):

thats 6.

OpenStudy (anonymous):

\[11*12= 132\] Being Canadian and working with centimeters mostly, I'm not entirely sure to answer yes or no, lemme just check real quick

OpenStudy (anonymous):

Yes, It is a reasonable size, i believe I don't really measure cows in my free time... :P

OpenStudy (anonymous):

lol thank you so much for your help

OpenStudy (anonymous):

No problem

OpenStudy (anonymous):

can you help me with a few more?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

I g2g here soon though just a couple

OpenStudy (anonymous):

The ratio of the lengths of corresponding sides of two similar polygons is 3:7 what is the ratio of their areas?

OpenStudy (anonymous):

details? lol

OpenStudy (anonymous):

thats all the question says

OpenStudy (anonymous):

Lolwut? can't answer anything from that :P

OpenStudy (anonymous):

yeah thats all the question says

OpenStudy (anonymous):

http://www.sonorahs.org/ourpages/auto/2010/4/28/54553484/M3C11CGD.pdf this website is the page im working on

OpenStudy (anonymous):

Sorry i got no clue w/o numbers

OpenStudy (anonymous):

ahh ok its the pictures

OpenStudy (anonymous):

well actual i need help with 5-14

OpenStudy (anonymous):

7:10 i think

OpenStudy (anonymous):

Eff i got no clue, i gotta peace out sorry couldn't stick around

OpenStudy (anonymous):

its okay thanks for your help

OpenStudy (anonymous):

No problem, just ask someone else they'll know hopefully and don't forget to thumb me ;)

OpenStudy (anonymous):

already did:)